# From Statics to Dynamics: Physics-Aware Image Editing with Latent Transition Priors

### v1

@@ 3.3 PhysicEdit Framework
Inspired by MetaQuery (pan2025transfer), we append $K$ (typically 64) learnable transition queries immediately after the reasoning prefix. These queries are processed by the frozen Qwen2.5-VL together with the text and source image, producing context-aware query embeddings. During training, we supervise these embeddings with a teacher branch derived from PhysicTran38K. We extract intermediate keyframes from the corresponding video, encode them using two complementary visual encoders (DINOv2 for semantic structure and the Qwen-Image-Edit VAE for fine-grained texture), and compress the resulting features into fixed-length latent features via feature extractors. We then map the transition-query embeddings through dual projection heads and align them to the structure- and texture-level targets, respectively. In this way, the transition queries learn to implicitly represent the missing evolution between $I_{src}$ and $I_{tgt}$, so that at inference time they can be instantiated solely from the source image, the instruction, and the physically-grounded reasoning. More details please refer to Appendix LABEL:sec:imple_details
@@
Since diffusion models exhibit coarse-to-fine generative trajectory, where global structure is formed before local texture is refined, we dynamically mix the compressed DINO latent feature $F_{\text{DINO}}$ and VAE latent feature $F_{\text{VAE}}$ according to the diffusion timestep. Let $t\in[0,1]$ denote the diffusion timestep, where larger $t$ corresponds to higher noise. We define the mixed transition feature as:
@@
When $t\rightarrow 1$, the modulation prioritizes $F_{\text{DINO}}$ to enforce a valid physical structure; when $t\rightarrow 0$, it shifts toward $F_{\text{VAE}}$ to refine texture details. The fused feature $F_{\text{tran}}^{(t)}$ is then injected into the diffusion model alongside the reasoning condition to guide the physics-aware generation.

### v2

# arXiv:2602.21778v2[cs.CV] 27 Feb 2026

@@ 3.3 PhysicEdit Framework
Inspired by MetaQuery (pan2025transfer), we append $K$ (typically 64) learnable transition queries immediately after the reasoning prefix. These queries are processed by the frozen Qwen2.5-VL together with the text and source image, producing context-aware query embeddings. During training, we supervise the query embeddings with pseudo-target transition features extracted from the associated video in PhysicTran38K. Specifically, we sample intermediate keyframes and encode them using two frozen visual encoders: DINOv2 for semantic structure and the Qwen-Image-Edit VAE for fine-grained texture. We then compress the encoder outputs into fixed-length pseudo-target embeddings, denoted as $F_{\text{DINO}}$ and $F_{\text{VAE}}$, using learnable feature extractors. In parallel, we map the transition-query embeddings through two projection heads to obtain the corresponding predictions, $\hat{F}_{\text{DINO}}$ and $\hat{F}_{\text{VAE}}$. We train the transition queries by aligning these predictions to the pseudo-targets. In this way, the transition queries learn to implicitly represent the missing evolution between $I_{src}$ and $I_{tgt}$, so that at inference time they can be instantiated solely from the source image, the instruction, and the physically-grounded reasoning. For more details, please refer to Appendix LABEL:sec:imple_details.
@@
Since diffusion models typically follow a coarse-to-fine trajectory, where global structure is formed before local texture is refined, we mix structure- and texture-level transition guidance according to the diffusion timestep. Let $t\in[0,1]$ denote the diffusion timestep, where larger $t$ corresponds to higher noise. During training, we construct the transition guidance from the video-derived pseudo-targets:
@@
When $t\rightarrow 1$, the guidance emphasizes $F_{\text{DINO}}$ to provide structurally consistent cues; when $t\rightarrow 0$, it emphasizes $F_{\text{VAE}}$ to refine texture-level details. $F_{\text{tran}}^{(t)}$ is then sent to MMDiT as condition. At inference time, the video-derived pseudo-targets are unavailable, and we instead instantiate the guidance from the transition queries’ predictions:
$\small\hat{F}_{\text{tran}}^{(t)}=t\cdot\hat{F}_{\text{DINO}}+(1-t)\cdot\hat{F}_{\text{VAE}}.$ (4)
The timestep-conditioned transition feature is injected into the diffusion model together with the physically-grounded reasoning condition to guide physics-aware generation.